# All-in-One Transferring Image Compression from Human Perception to Multi-Machine Perception

### v1

@@ 3.4. Optimization Strategy
(8) $\mathcal{L}_{\text{total}}=\lambda_{\text{rd}}\cdot\mathcal{L}_{rd}+\mathcal{L}_{\text{task}},$
@@
4. Experiments
Figure 3. Rate–accuracy performance comparison across different machine vision tasks between our method and other baselines. We report our method uses a single bitstream compare to other baseline generating a separate bitstream per task.
@@ 4. Experiments
Table 2. Comparison of rate–accuracy performance and trainable parameters (four tasks in total) across different tasks. BD-Rate and BD-Acc are computed with the base codec as anchor. Arrows indicate the preferred direction ($\downarrow$: lower is better, $\uparrow$: higher is better). (–) indicates BD-Rate is not available. $\Delta$m denotes the performance gain of our multi-task model over single-task counterparts (higher is better), where we report BD-acc as the task metric.  indicates that multi-task learning is not supported.
@@
BD-$\downarrow$ BD-$\uparrow$ BD-$\downarrow$ BD-$\uparrow$ BD-$\downarrow$ BD-$\uparrow$ BD-$\downarrow$ BD-$\uparrow$ Params (M) /$\Delta$m(%)
@@
TransTIC (Chen et al., 2023) -39.99% 4.36 -49.16% 2.74 -6.00% 0.23 -56.47% 0.73 6.48(21.4%)
AdaptICMH (Li et al., 2024b) -67.39% 8.45 -63.82% 4.78 -77.62% 2.76 -77.91% 1.09 1.15(3.7%)
Ours-Single Task -63.24% 9.04 -61.61% 4.45 -64.37% 2.86 -80.36% 1.08 0.55 (1.8%) -
Ours -74.89% 9.61 -78.91% 5.29 - 4.53 - 1.20 0.63 (2.1%) +0.24
4.1. Experimental Setup
@@
We conduct our experimental evaluation on the PASCAL-VOC dataset (Everingham et al., 2010), a widely-used benchmark in multi-task learning. Following prior works (Ye and Xu, 2022; Agiza et al., 2024; Vandenhende et al., 2020), we use the split of the PASCAL-Context subset with annotations for four vision tasks: semantic segmentation, human parsing, surface normal estimation, and saliency detection. The dataset contains 4,998 training images and 5,105 validation images. To meet codec constraints—which typically require image dimensions divisible by 64—all images are resized to $512\times 512$ during both training and testing. In our experiments, we jointly optimize these four tasks within our single unified framework.
@@
We use mean Intersection over Union (mIoU) to evaluate semantic segmentation, human part segmentation, and saliency prediction. For surface normal estimation, we report the root mean square error (RMSE) of the predicted angle differences. In addition, to assess the overall effectiveness of our multi-task learning, we adopt the multi-task performance metric $\Delta m$ introduced in (Maninis et al., 2019), which quantifies the average performance gain over its single-task counterparts. A higher $\Delta m$ indicates better multi-task generalization.
(9) $\Delta m=\frac{1}{T}\sum_{i=1}^{T}(-1)^{l_{i}}(M_{i}-M_{st,i})/M_{st,i}$
where $l_{i}=1$ if lower is better for task $i$, and $l_{i}=0$ otherwise. Here, $M_{i}$ and $M_{st,i}$ denote the performance of the multi-task and single-task models, respectively. In our experiments, we use BD-acc as the task performance metric $M_{i}$. BD-acc is detailed in Section 4.3.
4.2. Implementation Details
Our Framework
Our framework is built upon the pre-trained Lu2022-TIC base codec, which remains frozen during training. Trainable parameters are introduced via two lightweight modules: the task-agnostic Shared Adaptor (S-Adaptor) and the task-specific Task Adaptor (T-Adaptor). The S-Adaptor, shared across all tasks, is inserted into both encoder and decoder stages. In contrast, the T-Adaptor, placed only in the decoder stages, consists of four parallel paths corresponding to four downstream tasks. All S-Adaptors and T-Adaptor paths share the same architecture. The entire pipeline is trained end-to-end. Additionally, we analyze in the supplementary material how varying the reduction dimension within our adaptor modules impacts overall model performance.
Baselines
To benchmark our multi-task adaptation approach, we compare it with the following single-task tuning baselines: (1) Full Fine-tuning: All model parameters are fine-tuned independently for each task on top of the pre-trained base codec. (2) Adaptor Tuning (Li et al., 2024b): Several trainable lightweight adaptor modules are inserted into each encoder and decoder stage of the base codec. (3) Visual Prompt Tuning (VPT) (Chen et al., 2023): Trainable task-specific prompts are inserted into the transformer layers of the base codec.
All methods, including ours, are built upon the same base codec, Lu2022-TIC, and initialized with the pre-trained weights provided in (Chen et al., 2023) to ensure a fair comparison. For the Adaptor Tuning and Visual Prompt Tuning (VPT) baselines, we follow their official implementations and insert the PEFT modules into the encoder and decoder stages according to the configurations described in their respective papers.
@@ 4.2. Implementation Details
All experiments are conducted using PyTorch on a single NVIDIA RTX 4090 GPU. All models are trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. We apply a multi-step learning rate decay at epochs 2, 4, and 8 with a decay factor of 0.1. The total number of training epochs is set to 10.
We adopt the loss formulation described in Equation 8, combining rate-distortion loss and task objectives. For our method, task losses are balanced to a comparable scale following (Vandenhende et al., 2020), and the weights remain fixed across all experiments. We vary the coefficient $\lambda_{\text{rd}}$ within the range of 0.1 to 5.0 to explore the performance under different compression rates. The same strategy is applied to all baselines for fair comparison.
@@
For downstream task modeling, we adopted MTLoRA (Agiza et al., 2024), which utilizes a Swin-Tiny (Liu et al., 2021a) backbone coupled with a task-specific decoder head adapted from HRNet (Wang et al., 2020). The model is pre-trained on ImageNet-1K (Deng et al., 2009) and then fine-tuned on multiple downstream tasks using LoRA (Hu et al., 2021).
4.3. Main Results
We compare our method against VPT, Adaptor Tuning, Full Fine-tuning, and the base codec (denoted as TIC (Untuned)). fig. 3 presents rate–accuracy curves, showing the performance of each baseline under varying compression rates. Meanwhile, Following prior work (Chen et al., 2023), we compute BD-rate and BD-accuracy to quantitatively measure performance in Table 2 . BD-Rate calculate the average difference between two rate-distortion (RD) curves, indicating the bitrate savings at the same accuracy, while BD-acc compare the average accuracy difference between two rate-distortion (RD) curves, reflecting accuracy improvements at the same bitrate. For surface normal estimation (evaluated by RMSE, where lower is better), we negate the RMSE values before computing BD metrics.
@@ 4.3. Main Results
(1) Across all four downstream tasks and under varying compression rates, our method consistently outperforms all baselines. This demonstrates the strong generalization and task adaptability of our method.
(2) Our approach is significantly more parameter-efficient, equiring only 2.1% of the trainable parameters of the base codec—just half the size of the adaptor-based baseline.
@@
(4) Our architecture demonstrates beneficial task interactions. Our multi-task model achieves better performance than its single-task counterparts, consistently across tasks and bitrates, as also reflected in the positive $\Delta m$ score. This confirms the effectiveness of our multi-task architecture design.
4.4. Qualitive Results
We visualize task-specific reconstruction results in Figure 4. Overall, our model allocates more bitrate to texture-rich regions as expected. And the preserved frequency components are adaptively tailored to the needs of each downstream task. Specifically, region-level semantic segmentation tasks emphasize spatial coherence, such as object boundaries, shapes, and consistent textures. As a result, the model tends to retain more mid- and low-frequency components. Human part segmentation, being even more semantically constrained, exhibits a stronger focus on low-frequency structures. In contrast, saliency detection aims to highlight visually distinctive regions, such as object edges, fine textures, and sharp intensity contrasts, which are typically characterized by high-frequency information. Therefore, reconstructions optimized for saliency retain more high-frequency details. Surface normals estimation, which predicts the per-pixel orientation of surfaces in 3D space, relies on both mid-frequency cues for smooth geometric transitions and certain high-frequency signals to capture fine-grained surface details. Consequently, its frequency distribution lies between that of segmentation and saliency tasks. We provide more visual reconstruction results in the supplementary materials.
Figure 4. Our Reconstruction results at 0.1168 bpp. The first column shows reconstructions optimized for different vision tasks. The second column displays the corresponding downstream task predictions. The third column presents task-specific decoding output residuals, to better visualization, we generate residual heatmaps that highlight task-relevant differences. The fourth column visualizes the frequency-domain transformation of the reconstructed images. The fifth column includes the ground-truth images, corresponding frequency-domain representations, and the bit allocation maps after being compressed.
[…]
main_results_visual
4.5. Ablation Study
Figure 5. Comparison of multi-scale fusion strategies: (a) Sum Fusion, (b) Cross-Scale Attention, and (c) Gate Fusion.
[…]
Figure 6. Comparison of multi-scale fusion strategies: (a) Sum Fusion, (b) Cross-Scale Attention, and (c) Gate Fusion.
[…]
Multi-Scale Fusion Strategy
Our T-Adaptors at different decoder stages generate multi-scale features that capture varying levels of information from the input image. Shallow-stage features, with smaller spatial resolutions, tend to represent higher-level semantics and geometric structures, while deeper-stage features capture finer details such as textures and edges. Although all scales are essential for accurate image reconstruction, they contribute differently to each downstream task. Effectively aggregating these multi-scale features is crucial for robust task performance. To this end, we compare three multi-scale fusion strategies, illustrated in Figure 6: (a) Sum Fusion: Features from different scales are spatially aligned and fused via element-wise summation. (b) Cross-Scale Attention: Attention modules are applied across scales to adaptively weigh the importance of each feature map. (c) Gate Fusion: Learnable gates are used to control the contribution of each scale, enabling dynamic feature selection conditioned on task requirements.
We summarize the rate–accuracy comparison of different fusion strategies in Figure 5. Both Gate Fusion and Cross-Scale Attention achieve comparable and superior performance to Sum Fusion. Gate Fusion achieves slightly better results at lower bitrates, while Cross-Scale Attention excels under higher bitrates, which intuitively due to its global modeling capability benefiting from dense features. Considering the trade-off between performance, parameter count, and computational complexity, we adopt Gate Fusion in our main results reported in Figure 3 and Table 2.
@@ 4.5. Ablation Study
Table 3. BD-rate and BD-acc Performance of Semantic Segmentation (SemSeg) and Human Parsing (Human) comparison unber varing task number (set SemSeg as anchor task).
@@
+Human+Sal -70.13%/9.39 -74.85%/5.22
+Human+Sal+Normals -74.89% /9.61 -78.91%/5.29
@@
Although our method requires the fewest trainable parameters to adapt the base codec, the parallel multi-path design of the Task Adaptors introduces additional architectural complexity, which may affect computational efficiency. To assess this impact, we follow prior works (Chen et al., 2023) and compare the complexity of competing methods using kilo multiply–accumulate operations per pixel (kMACs/pixel).
@@
Table 4. Comparison of the kMACs/pixel and model size.
@@
TIC (Full-Finetune) 142.5 188.5 7.51 $\times$ 4
TransTIC (Chen et al., 2023) 332.0 202.6 1.62 $\times$ 4
AdaptICMH (Li et al., 2024b) 157.2 203.2 0.29 $\times$ 4
@@
5. Conclusion
@@ Conclusion
In conclusion, we introduce a novel asymmetric adaptation framework that enables efficient transfer of a human-centric base codec to a multi-task vision codec. Specifically, we design two lightweight plug-in modules: a task-agnostic Shared Adaptor for generalizable feature sharing, and a task-specific Task Adaptor for decoupled task modulation. The entire framework follows the PEFT training paradigm and requires only 2.1% of the trainable parameters to transfer to four downstream tasks, demonstrating its effectiveness under task number scaling.
Overall, this work offers a practical and effective solution for machine-centric image compression, and to the best of our knowledge, represents the first attempt at multi-task transfer in the codec transfer domain.

### v2

@@ 3.5 Optimization Strategy
$\mathcal{L}_{\text{total}}=\lambda_{\text{rd}}\cdot\mathcal{L}_{rd}+\mathcal{L}_{\text{task}},$ (9)
@@
4 Experiments
Figure 2: Rate–Accuracy performance comparison across four machine vision tasks on PASCAL-Context.
@@ 4 Experiments
TransTIC -39.99% 4.36 -49.16% 2.74 -6.00% 0.23 -56.47% 0.73 6.48(21.4%)
AdaptICMH -67.39% 8.45 -63.82% 4.78 -77.62% 2.76 -77.91% 1.09 1.15(3.7%)
Ours -74.89% 9.61 -78.91% 5.29 - 4.53 - 1.20 0.63 (2.1%) +0.25
Table 2: Comparison of Rate–Accuracy performance and trainable parameters across four tasks on PASCAL-Context. BD-Rate and BD-Acc are calculated with the base codec as anchor. Arrows indicate the preferred direction ($\downarrow$: lower is better, $\uparrow$: higher is better). (–) indicates unavailableBD-Rate. $\Delta$m represents the BD-Acc improvement of our multi-task model over the single-task baseline (AdaptICMH). ✗ indicates that multi-task learning is not supported.
Figure 3: Rate–Accuracy performance comparison across different machine vision tasks on NYUD-V2.
[…]
Full Fine-Tuning -67.45% 3.61 -69.10% 0.16 -21.80% 0.51 30.04(100%)
TransTIC -10.82% 0.57 -45.00% 0.11 -25.79% 0.52 4.86(16.18%)
AdaptICMH -60.42% 3.24 -58.97% 0.16 18.50% 0.76 0.86(2.9%)
Ours -66.06% 4.06 -79.67% 0.18 -46.70% 1.11 0.50 (1.7%) +0.28
Table 3: Comparison of Rate–Accuracy performance and trainable parameters across three tasks on NYUD-V2.
4.1 Experimental Setup
@@
We conduct our experimental evaluation on two widely used benchmarks in multi-task learning: PASCAL-Context [14] and NYUD-V2 [37]. On PASCAL-Context, we evaluate four tasks: Semantic Segmentation, Human Parsing, Surface Normal Estimation, and Saliency Detection. On NYUD-V2, we consider three tasks: semantic segmentation, depth estimation, Surface Normals Estimation. To meet codec constraints—which typically require image dimensions divisible by 64—all images are resized to $\times 512$ 512 during both training and testing.
@@
We use mean Intersection over Union (mIoU) to evaluate semantic segmentation, human parsing, and saliency detection. For depth estimation and surface normal estimation, we report Root Mean Square Error (RMSE) of the predicted depth and angle differences. In addition, to assess the overall effectiveness of our multi-task learning framework, we adopt the multi-task performance metric $\Delta m$ introduced in [33], which quantifies the average performance gain over the corresponding single-task baselines. A higher $\Delta m$ indicates better multi-task generalization.
$\Delta m=\frac{1}{T}\sum_{i=1}^{T}(-1)^{l_{i}}(M_{i}-M_{st,i})/M_{st,i}$ (10)
where $l_{i}=1$ if lower is better for task $i$, and $l_{i}=0$ otherwise. Here, $M_{i}$ and $M_{st,i}$ denote the performance of the multi-task and single-task models, respectively. In our experiments, we use BD-Acc as the task performance metric $M_{i}$, AdaptICMH the single-task baseline.
4.2 Implementation Details
The training of Our framework strictly follows the coding-for-machine paradigm. Our method outputs a decoded image for each task (with one single inference), which is then processed by the downstream model. The loss is computed between downstream model’s predictions and its ground-truth labels, and the final training objective is a weighted sum of task-specific losses, with weights adopted from [39]. Gradient flow follows the parameter-efficient fine-tuning (PEFT) pipeline: gradients are propagated from the downstream models back through the compression model, including both the base codec and our tuning modules. Importantly, thanks to our parallel adaptor design, the gradients of each task-specific adaptor are primarily attributed to its corresponding task-specific loss, which equips the model with the capacity for task decoupling. Additionally, to benchmark our multi-task adaptation approach, we compare it with the following single-task tuning baselines: (1) Full Fine-tuning: All model parameters are fine-tuned independently for each task on top of the pre-trained base codec. (2) Adaptor Tuning [24]: Several trainable lightweight adaptor modules are inserted into each encoder and decoder stage of the base codec. (3) Visual Prompt Tuning (VPT) [8]: Trainable task-specific prompts are inserted into the transformer layers of the base codec.
All methods, including ours, are built upon the same base codec, Lu2022-TIC, and initialized with the pre-trained weights provided in [8] to ensure a fair comparison. For the Adaptor Tuning and VPT baselines, we follow their official implementations and insert the PEFT modules into the encoder and decoder stages according to the configurations described in their respective papers.
@@ 4.2 Implementation Details
All experiments are conducted using PyTorch on a single NVIDIA RTX 4090 GPU. Models are trained with the Adam optimizer, starting from an initial learning rate of $1\times 10^{-4}$, and employing a multi-step learning rate decay at epochs 2, 4, and 8 with a decay factor of 0.1. The total number of training epochs is set to 10. We adopt the loss formulation described in Equation 9, which combines the rate-distortion loss with task-specific objectives. To explore performance under different compression rates, we vary the rate-distortion trade-off coefficient $\lambda_{\text{rd}}$ in the range of 0.1 to 5.0.
@@
For downstream task modeling, we adopted MTLoRA [1], which utilizes a Swin-Tiny [31] backbone coupled with a task-specific decoder head adapted from HRNet [40]. The model is pre-trained on ImageNet-1K [13] and then fine-tuned on multiple downstream tasks using task-specific LoRA [19] modules.
4.3 Main Results
We compare our method against Full Fine-tuning, Adaptor Tuning, VPT, and the base codec (denoted as TIC (Untuned)). Figure 2 and Figure 3 show the Rate–Accuracy curves on PASCAL-Context and NYUD-V2, respectively, illustrating the performance of each baseline under varying compression rates. Following prior work [8], we additionally report BD-Rate and BD-Accuracy in Table 2 and Table 3 to provide a quantitative comparison. BD-Rate measures the average bitrate savings at equivalent accuracy, while BD-Accuracy reflects the average accuracy improvement at equivalent bitrate. Additional qualitative results are provided in the appendix to further illustrate the visual performance of our method.
@@ 4.3 Main Results
(1) Across all four/three downstream tasks and under varying compression rates, our method consistently outperforms all baselines. This demonstrates the strong generalization and task adaptability of our method.
(2) Our approach is rather more parameter-efficient, equiring only 2.1%/1.7% of the trainable parameters of the base codec—just half the size of the adaptor-based baseline.
@@
(4) Our architecture demonstrates beneficial task interactions. Our multi-task model achieves better performance than its single-task baseline, consistently across tasks and bitrates, as also reflected in the positive $\Delta m$ score.
4.4 Ablation Study
@@
+Human/Sal -70.13%/9.39 -74.85%/5.22
+Human/Sal/Normal -74.89% /9.61 -78.91%/5.29
Table 4: BD-Rate and BD-Acc Performance of Semantic Segmentation and Human Parsing comparison under varing task number (set SemSeg as anchor task).
@@
Although our method requires the fewest trainable parameters to adapt the base codec, the parallel multi-path design of the Task Adaptors introduces additional architectural complexity, which may affect computational efficiency. To assess this impact, we follow prior works [8] and compare the complexity of competing methods using kilo multiply–accumulate operations per pixel (kMACs/pixel).
@@
Full-Finetune 142.5 188.5 7.51 $\times$ 4
TransTIC 332.0 202.6 1.62 $\times$ 4
AdaptICMH 157.2 203.2 0.29 $\times$ 4
@@
Table 5: Comparison of the kMACs/pixel and model size.
Sparse Prediction Task
Owing to the reliance of our setting on existing multi-task learning benchmarks, most evaluations are limited to dense prediction tasks. While the tasks we included could already reflect the model’s ability to capture both global and local information—semantic segmentation, for instance, depends on a global receptive field (e.g., locating a person in the scene), whereas depth and surface normal estimation rely more on local receptive fields to capture fine-grained edge and texture details—we are further interested in understanding how well our framework generalizes to structurally different tasks.
To this end, we extend our evaluation to sparse prediction tasks using the COCO 2017 [25], focusing on two representative tasks: object detection and instance segmentation. As shown in Figure 4, our method achieves performance that is comparable to—or even exceeds—that of single-task baselines, demonstrating its robustness and generalizability beyond the dense prediction setting.
Figure 4: Rate–Accuracy performance comparison of Object Detection and Instance Segmentation on COCO-2017.
5 Conclusion
@@ Conclusion
In conclusion, we propose a novel multi-task adaptation framework that enables efficient transfer of a human-centric base codec to a multi-task vision codec. Specifically, we introduce an asymmetric adaptor tuning architecture, which consists of a task-agnostic, single-path encoder adaptor and task-specific, parallel-path decoder adaptors. On top of this structure, we design two feature propagation modules to facilitate both inter-task and inter-scale information exchange. We validate the effectiveness of our framework on multiple benchmark datasets.
Overall, this work presents a practical and effective solution aligned with the coding-for-machine paradigm, and to the best of our knowledge, constitutes the first attempt to enable multi-task transfer in the context of codec adaptation.